how many metric tons of chromiun must be added to 2.50 metric tons of stainless steel to raise the percent chromium from 11% to 18%
Answer by longjonsilver(2297)

questions like these are always about "BEFORE" + "ADDED" = "AFTER"

BEFORE:
2.5*0.11 of 11% chromium steel
x*1.00 of 100% chromium --> this is the "tricky" part, realising that the chromium is 100%

AFTER:
(2.5+x)*0.18

so, these 2 are equivalent:
2.5*0.11 + x = (2.5+x)*0.18

0.275 + x = 0.45 + 0.18x
0.275 + 0.82x = 0.45
0.82x = 0.175
--> x = 0.21 metric tonnes

jon.